y=(-1/2) x^2-3x+11
find the max or min value ofy and the value of x when it occurs. thanks.

Answer by stanbon(75887)

y=(-1/2) x^2-3x+11
find the max or min value of y and the value of x when it occurs. thanks.
----
Since the coefficient of x^2 is negative, y has a maximum value.
--------------------
Note:: Max occurs when x = -b/(2a) = 3/(2(-1/2)) = 3/-1 = -3
------
The max value is f(-3) = (-1/2)(9)-3(-3)+11 = -4.5 + 20 = 15.5
----------------
Cheers,
Stan H.
----------------